OpenStudy (anonymous):

Compare and contrast the two quadratic equations below. In order to receive full credit, use complete sentences to describe the following: The direction each parabola opens The vertex of each parabola y = x2 − 2x y = −2x2 + 4x − 3

OpenStudy (anonymous):

How would I do this?

mathslover (mathslover):

Well the direction will be opening upwards as \(x^2\) term is positive

OpenStudy (anonymous):

what equation forms for parabolas do you know?

OpenStudy (anonymous):

you could graph them both...

mathslover (mathslover):

Case 1 : \(y= ax^2 + bx + c\) when \(a>0\) , the parabola is opening upwards Case 2 : \(y = ax^2 + bx + c\) when \(a<0\) , the parabola is opening downwards.

OpenStudy (anonymous):

Ok, thanks . . .

mathslover (mathslover):

Differentiate : \(y = x^2-4x\) , what do you get?

OpenStudy (anonymous):

ummmm . . . y = (x-4) x ?

mathslover (mathslover):

Do you know differentiation?

OpenStudy (anonymous):

apparently not

OpenStudy (anonymous):

it will come later

OpenStudy (anonymous):

Ok . . . I am officially confused.

mathslover (mathslover):

Hmn. Let us do it with other method then!

OpenStudy (anonymous):

Ok fine by me . . .

mathslover (mathslover):

Well is it \(x^2-2x\) or \(x^2-4x\) ?

OpenStudy (anonymous):

The 1st one why wouldn't it be?

mathslover (mathslover):

Hmn just asking as if it was \(x^2-4x\) then it may have been little bit easier for me.

OpenStudy (anonymous):

The is 2 grades above me . . .

mathslover (mathslover):

ok1 no problem .. Let us do it as \(x^2-2x + 2 - 2 = y\) is it fine for you?

OpenStudy (anonymous):

Yes

mathslover (mathslover):

Ok fine, forget it. I have to think more.

OpenStudy (anonymous):

Confused . . .

mathslover (mathslover):

Wait for 1 min. more frank

OpenStudy (anonymous):

Ok

mathslover (mathslover):

Actually the problem I am having is that I have to complete the square of \(x^2-2x\) ,,,

mathslover (mathslover):

Ok got it.

mathslover (mathslover):

\(x^2- 2x +1-1 = y\) Fine again?

OpenStudy (anonymous):

Yes

mathslover (mathslover):

Now Can I write \(x^2-2x+1 = x^2 - x - x +1 = x(x-1) - (x-1) = (x-1)(x-1)\) ?

mathslover (mathslover):

Or \(x^2-2x +1 = (x-1)^2\) ?

OpenStudy (anonymous):

Bleg . . . I don't realy get any of this . . .

mathslover (mathslover):

hmn : what is (a-b)^2?

OpenStudy (anonymous):

(a-b)(a-b)?

mathslover (mathslover):

Expand that frank.

OpenStudy (anonymous):

If you have \[(a-b)^{2} \] wouldn't you just be multiplying them by themselves?

mathslover (mathslover):

\((a-b)(a-b)\) = \(a(a) - a(b) -b(a) -b(-b)\) right?

OpenStudy (anonymous):

I think so . . .

mathslover (mathslover):

now try to simplify it: \(a(a) - a(b) - b(a) -b(-b)\)